# Tackling Decision Processes with Non-Cumulative Objectives using Reinforcement Learning

### v1

@@ Abstract
Markov decision processes (MDPs) are used to model a wide variety of applications ranging from game playing over robotics to finance. Their optimal policy typically maximizes the expected sum of rewards given at each step of the decision process. However, a large class of problems does not fit straightforwardly into this framework: Non-cumulative Markov decision processes (NCMDPs), where instead of the expected sum of rewards, the expected value of an arbitrary function of the rewards is maximized. Example functions include the maximum of the rewards or their mean divided by their standard deviation. In this work, we introduce a general mapping of NCMDPs to standard MDPs. This allows all techniques developed to find optimal policies for MDPs, such as reinforcement learning or dynamic programming, to be directly applied to the larger class of NCMDPs. Focusing on reinforcement learning, we show applications in a diverse set of tasks, including classical control, portfolio optimization in finance, and discrete optimization problems. Given our approach, we can improve both final performance and training time compared to relying on standard MDPs.
@@ 1 Introduction
Markov decision processes (MDPs) are used to model a wide range of applications where an agent iteratively interacts with an environment during a trajectory. Important examples include robotic control [Andrychowicz et al. (2020a)], game playing [Mnih et al. (2015)], or discovering algorithms [Mankowitz et al. (2023)]. At each time step $t$ of the MDP, the agent chooses an action based on the state of the environment and receives a reward $r_{t}$. The agent’s goal is to follow an ideal policy that maximizes
@@
where $T$ is the length of the trajectory and the expectation value is taken over both the agent’s probabilistic policy $\pi$ and the probabilistic environment. There exist a multitude of established strategies for finding (approximately) ideal policies of MDPs, such as dynamic programming and reinforcement learning [Sutton & Barto (2018)].
A limitation of the framework of MDPs is the restriction to ideal policies that maximize Equation 1, while a large class of problems cannot straightforwardly be formulated this way. For example, in weakest-link problems, the goal is to maximize the minimum rather than the sum of rewards, e.g. in network routing one wants to maximize the minimum bandwidths along a path [Cui & Yu (2023)]. In finance, the Sharpe ratio, i.e. the mean divided by the standard deviation of portfolio gains, is an important figure of merit of an investment strategy, since maximizing it will yield more risk-averse strategies than maximizing the sum of portfolio gains [Sharpe (1966)]. We therefore require a method to tackle non-cumulative Markov decision processes (NCMDPs), where instead of the expected sum of rewards, the expectation value of an arbitrary function of the rewards is maximized. First steps in this direction have already been taken, but are currently limited to specific settings, e.g. by being restricted to only certain MDP solvers, deterministic environments, and restricted classes of non-cumulative objectives (for more details on these approaches see Section 4). The main contribution of this paper is twofold:
  • First, we provide a theoretical framework for a general and easy-to-implement mapping of NCMDPs to corresponding standard MDPs. This allows the direct application of advanced MDP solvers such as reinforcement learning or dynamic programming to tackle also NCMDPs (see Figure 1).
  • Next, we perform numerical experiments focusing on reinforcement learning. We show applications in classical control problems, portfolio optimization, and discrete optimization problems using e.g. the non-cumulative $\max$ objective and the Sharpe ratio. Using our framework, we find improvements in both training time and final performance as compared to relying on standard MDPs.
@@ 2 Theoretical Analysis
Non-cumulative Markov decision processes
@@
NCMDPs still have a Markovian transition probability distribution but their return, i.e. Equation 2, depends on the rewards in a non-cumulative and therefore non-Markovian way. Thus, NCMDPs are a generalization of MDPs. Due to this distinction, MDP solvers cannot straightforwardly be used for NCMDPs. To solve this problem, we describe a general mapping from an NCMDP $\tilde{M}$ with rewards $\tilde{r}_{t}$ and states $\tilde{s}_{t}$ to a corresponding MDP $M$ with adapted states $s_{t}$ and adapted rewards $r_{t}$ but the same actions. The mapping is chosen such that the optimal policy of $\tilde{M}$ is equivalent to the optimal policy of $M$. Therefore, a solution for $\tilde{M}$ can readily be obtained by solving $M$ with existing MDP solvers, as shown in Figure 1.
As a first step, we define
$r_{t}=f(\tilde{r}_{0},\dots,\tilde{r}_{t})-f(\tilde{r}_{0},\dots,\tilde{r}_{t-1}).$ (3)
As $\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}r_{t}\right]=\mathbb{E}_{\pi}\left[f(\tilde{r}_{0},\dots,\tilde{r}_{T-1})\right]$, the ideal policy of $M$ is also the ideal policy of $\tilde{M}$. However, the $r_{t}$ are now non-Markovian, since they depend on rewards in the past of the trajectory. Recovering theoretical convergence guarantees available for MDP solvers requires adapting our state space to recover the Markov property of $M$. In principle, methods developed based on past [Bacchus et al. (1996, 1997)] or future [Thiébaux et al. (2006)] linear temporal logic could be employed to achieve this. However, we argue that for a wide range of functions $f$, a more straightforward approach is possible: The primary challenge is that we need to access the non-cumulative return in a purely Markovian manner. This can be achieved by extending the state in a way that preserves all the information necessary to compute this return. The additional state information $x_{t}$ needs to be sufficient to compute $r_{t}$, i.e. the temporal difference of the non-cumulative returns $f$. Concretely, we need to define a vector $x_{t}$ and two functions $y$ and $z$, such that the following relations hold:
$\begin{gathered}s_{t}=(\tilde{s}_{t},x_{t})\text{ with }x_{t+1}=y(x_{t},\tilde{r}_{t}),\\ r_{t}=f(\tilde{r}_{0},\dots,\tilde{r}_{t})-f(\tilde{r}_{0},\dots,\tilde{r}_{t-1})=z(x_{t},\tilde{r}_{t}).\end{gathered}$ (4)
With the Dirac delta function $\delta$, the transition probabilities $p$ of $M$ are then given in terms of the transition probabilities $\tilde{p}$ of $\tilde{M}$ by
$p(r_{t},s_{t+1}|s_{t},a_{t})=\tilde{p}(\tilde{r}_{t},\tilde{s}_{t+1}|\tilde{s}_{t},a_{t})\delta\left[x_{t+1}-y(x_{t},\tilde{r}_{t})\right]\delta\left[r_{t}-z(x_{t},\tilde{r}_{t})\right].$ (5)
There are several possible choices for $x_{t}$, $y$, and $z$. For example, it is always possible to choose $x_{t}=(r_{0},\dots,r_{t-1})$. However, this is undesirable because it leads to a state size that grows with the trajectory length. Fortunately, it is possible to find $x_{t}$ of a small, constant size for many functions $f$. For example, for $f(\tilde{r}_{0},\dots,\tilde{r}_{t})=\min(\tilde{r}_{0},\dots,\tilde{r}_{t})$, we can choose $r_{0}=x_{1}=\tilde{r}_{0}$ followed by $x_{t}=y(x_{t},\tilde{r}_{t})=\min(x_{t},\tilde{r}_{t})$, and $r_{t}=z(x_{t},\tilde{r}_{t})=\min(0,\tilde{r}_{t}-x_{t})$. More examples are shown in Table 1. By mapping the NCMDP to an MDP, all guarantees available for MDP solvers, e.g. convergence proofs for Q-learning [Watkins & Dayan (1992)], or the policy gradient theorem for policy-based methods [Sutton et al. (1999)], directly apply also to NCMDPs.
@@
$\bm{f(\tilde{r}_{0}},\dots,\bm{\tilde{r}_{T-1})}$ State adaption $\bm{x_{t}}$ Adapted reward $\bm{r_{t}}$
$\max(\tilde{r}_{0},\dots,\tilde{r}_{T-1})$ $x_{1}=\tilde{r}_{0}$, $x_{t+1}=\max(x_{t},\tilde{r}_{t})$ $r_{t}$ = $\max(0,\tilde{r}_{t}-x_{t})$
$\min(\tilde{r}_{0},\dots,\tilde{r}_{T-1})$ $x_{1}=\tilde{r}_{0}$, $x_{t+1}=\min(x_{t},\tilde{r}_{t})$ $r_{t}$ = $\min(0,\tilde{r}_{t}-x_{t})$
Sharpe ratio $\frac{\mathrm{MEAN}(\tilde{r}_{0},\dots,\tilde{r}_{T-1})}{\mathrm{STD}(\tilde{r}_{0},\dots,\tilde{r}_{T-1})}$ $x_{0}=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}$, $x_{t+1}=\begin{bmatrix}\frac{x_{t}^{(2)}}{x_{t}^{(2)}+1}x_{t}^{(0)}+\frac{1}{x_{t}^{(2)}+1}\tilde{r}_{t}\\ \frac{x_{t}^{(2)}}{x_{t}^{(2)}+1}x_{t}^{(1)}+\frac{1}{x_{t}^{(2)}+1}\tilde{r}_{t}^{2}\\ \ x_{t}^{(2)}+1\end{bmatrix}$
$r_{t}=\frac{x_{t+1}^{(0)}}{\sqrt{x_{t+1}^{(1)}-x_{t+1}^{(0)^{2}}}}$
$-\frac{x_{t}^{(0)}}{\sqrt{x_{t}^{(1)}-x_{t}^{(0)^{2}}}}$
$\underset{k\in[-1,T-1]}{\max}\sum_{t=0}^{k}\tilde{r}_{t}$ $x_{0}=0$, $x_{t+1}=\max(0,x_{t}-\tilde{r}_{t})$ $r_{t}$ = $\max(0,\tilde{r}_{t}-x_{t})$
$\tilde{r}_{0}\tilde{r}_{1}\dots\tilde{r}_{T-1}$ $x_{0}=1$, $x_{t+1}=\tilde{r}_{t}x_{t}$ $r_{t}=x_{t+1}-x_{t}$
Harmonic mean $\frac{1}{\frac{1}{\tilde{r}_{0}}+\dots+\frac{1}{\tilde{r}_{T-1}}}$ $x_{0}=0,x_{t+1}=x_{t}+\frac{1}{\tilde{r}_{t}}$ $r_{t}=\frac{1}{x_{t+1}}-\frac{1}{x_{t}}$
$\delta^{T}\sum_{t=0}^{T-1}\tilde{r}_{t}$,
$\delta\in(0,1)$
$x_{0}=\begin{bmatrix}0\\ 0\end{bmatrix}$, $x_{t+1}=\begin{bmatrix}x_{t}^{(0)}+\tilde{r}_{t}\\ x_{t}^{(1)}+1\end{bmatrix}$
$r_{t}=\delta^{x_{t+1}^{(1)}}x_{t+1}^{(0)}$
$-\delta^{x_{t}^{(1)}}x_{t}^{(0)}$
@@
We demonstrate the need for the additional state information $x_{t}$ in the two-step NCMDP depicted in Figure 2. Here, the goal is to maximize the expected minimum of the rewards, i.e. $\min(\tilde{r}_{0},\tilde{r}_{1})$. If $\tilde{r}_{0}=1$ in the first step, the ideal policy is to choose $a_{1}$ in the second step, while if $\tilde{r}_{0}=-1$ the agent should choose $a_{0}$. Therefore, the choice of the ideal action depends on information contained in the past rewards, which is captured by $x_{1}=\tilde{r}_{0}$. We show the MDP constructed from the NCMDP as described above in the Appendix in Figure A1 and its state-action value function found by value iteration in Table A1.

### v3

@@ Abstract
Markov decision processes (MDPs) are used to model a wide variety of applications ranging from game playing over robotics to finance. Their optimal policy typically maximizes the expected sum of rewards given at each step of the decision process. However, many real-world problems do not fit straightforwardly into this framework: Non-cumulative Markov decision processes (NCMDPs), where instead of the expected sum of rewards, the expected value of an arbitrary function of the rewards is maximized. Example functions include the maximum of the rewards or their mean divided by their standard deviation. In this work, we introduce a general mapping of NCMDPs to standard MDPs. This allows all techniques developed to find optimal policies for MDPs, such as reinforcement learning or dynamic programming, to be directly applied to the larger class of NCMDPs. We demonstrate the effectiveness of our approach in diverse reinforcement learning tasks, including classical control, financial portfolio optimization, and discrete optimization. Our approach improves both final performance and training efficiency compared to relying on standard MDPs.
Keywords: Markov Decision Processes, Reinforcement Learning, Deep Learning, Discrete Optimization, Portfolio Optimization
@@ 1 Introduction
Markov decision processes (MDPs) are used to model a wide range of applications where an agent iteratively interacts with an environment during a trajectory. Important examples include robotic control (Andrychowicz et al., 2020a), game playing (Mnih et al., 2015), or discovering algorithms (Mankowitz et al., 2023). At each time step $t$ of the MDP, the agent chooses an action based on the state of the environment and receives a reward $r_{t}$. The agent’s goal is to follow an ideal policy that maximizes
@@
where $T$ is the length of the trajectory and the expectation value is taken over both the agent’s probabilistic policy $\pi$ and the probabilistic environment. There exist a multitude of established strategies for finding (approximately) ideal policies of MDPs, such as dynamic programming and reinforcement learning (Sutton and Barto, 2018).
A limitation of the framework of MDPs is the restriction to ideal policies that maximize Equation 1, while a large class of problems cannot straightforwardly be formulated this way. For example, in weakest-link problems, the goal is to maximize the minimum rather than the sum of rewards, e.g. in network routing one wants to maximize the minimum bandwidths along a path (Cui and Yu, 2023). In finance, the Sharpe ratio, i.e. the mean divided by the standard deviation of portfolio gains, is an important figure of merit of an investment strategy, since maximizing it will yield more risk-averse strategies than maximizing the sum of portfolio gains (Sharpe, 1966). We therefore require a method to tackle non-cumulative Markov decision processes (NCMDPs), where instead of the expected sum of rewards, the expectation value of an arbitrary function of the rewards is maximized. The main contribution of this paper is twofold:
  1. 1.
    We expand the scope of reinforcement learning by introducing a theoretical framework that maps NCMDPs to standard MDPs. This mapping enables direct application of advanced MDP solvers, including reinforcement learning and dynamic programming, to NCMDPs without modification. Our approach accommodates general non-cumulative objectives and is straightforward to implement.
    Context: Existing methods for solving NCMDPs are restricted to either maximizing the highest reward (Quah and Quek, 2006; Veviurko et al., 2024) or handling a limited set of objectives using Q-learning in deterministic settings (Cui and Yu, 2023). In contrast, our framework generalizes to arbitrary non-cumulative objectives in both deterministic and stochastic environments, albeit at the cost of increased state space size. In practice, this is not a fundamental limitation for deep reinforcement learning techniques, since they are well equipped to handle even exponentially large state spaces. Moreover, unlike prior approaches that require modifying the reinforcement learning algorithm, our method treats both the environment and learning algorithm as black boxes, simplifying implementation.
  2. 2.
    We conduct numerical experiments demonstrating the effectiveness of deep reinforcement learning on NCMDPs across complex applications with large state and action spaces, including portfolio optimization, classical control, and discrete optimization. Our results indicate improvements in training efficiency and final performance compared to prior reinforcement learning approaches.
    Context: The problems we investigate involve non-cumulative objectives and stochastic environments that place them beyond the scope of previous NCMDP-solving methods. As a result, prior state-of-the-art reinforcement learning approaches were constrained by the need to approximate non-cumulative objectives using standard MDPs. Our method overcomes this limitation by allowing direct specification of the exact objective to the reinforcement learning agent, leading to improved performance.
@@ 2.1 Preliminaries
2.2 Non-cumulative Markov Decision Processes
@@
To accommodate trajectories of arbitrary finite length, we require $f$ to be a family of functions consisting of a function $f_{t}:\mathbb{R}^{t}\rightarrow\mathbb{R}$ for each $t\in\mathbb{N}$. For brevity, we denote $f_{t}(\tilde{r}_{0},\tilde{r}_{1},\dots,\tilde{r}_{t-1})=f(\tilde{r}_{0},\tilde{r}_{1},\dots,\tilde{r}_{t-1})$. NCMDPs still have a Markovian transition probability distribution but their return, i.e. Equation 2, can depend on the rewards in a non-cumulative and therefore non-Markovian way. Thus, NCMDPs are a generalization of MDPs. Due to this distinction, MDP solvers cannot straightforwardly be used for NCMDPs. To solve this problem, we describe a general mapping from an NCMDP $\tilde{M}$ to a corresponding MDP $M$ with adapted states and adapted rewards but the same actions. The mapping is chosen such that the optimal policy of $\tilde{M}$ is equivalent to the optimal policy of $M$. Therefore, a solution for $\tilde{M}$ can readily be obtained by solving $M$ with existing MDP solvers, as shown in Figure 1. In the following, we formalize these ideas.
Definition 1
Given an NCMDP $\tilde{M}$ with rewards $\tilde{r}_{t}$, states $\tilde{s}_{t}$, actions $a_{t}$, the Markovian state transition probability distribution $\tilde{p}(\tilde{r}_{t},\tilde{s}_{t+1}|\tilde{s}_{t},a_{t})$, and non-cumulative objective function $f$, we define a corresponding MDP $M$ with rewards $r_{t}$, states $s_{t}$, the same actions $a_{t}$, and state transition probability distribution $p(r_{t},s_{t+1}|s_{t},a_{t})$ with functions $\rho$ and $u$, and vectors $h_{t}$ with initial values $h_{0}$ such that
$r_{t}=f(\tilde{r}_{0},\dots,\tilde{r}_{t})-f(\tilde{r}_{0},\dots,\tilde{r}_{t-1})=\rho(h_{t},\tilde{r}_{t}),$ (3)
$s_{t}=(\tilde{s}_{t},h_{t})\text{ with }h_{t+1}=u(h_{t},\tilde{r}_{t}),$ (4)
$p\big{(}r_{t},s_{t+1}{=}(\tilde{s}_{t+1},h_{t+1})|s_{t}{=}(\tilde{s}_{t},h_{t}),a_{t}\big{)}=\sum_{\tilde{r}_{t}}\tilde{p}(\tilde{r}_{t},\tilde{s}_{t+1}|\tilde{s}_{t},a_{t})\delta_{h_{t+1},u(h_{t},\tilde{r}_{t})}\delta_{r_{t},\rho(h_{t},\tilde{r}_{t})},$ (5)
where $\delta$ is the Kronecker delta. For continuous probability distributions, the sum should be replaced by an integral and the $\delta$ by Dirac delta functions.
We now provide some intuition for this definition: Equation 3 ensures that the return of $M$, i.e. $\sum_{t=0}^{T-1}r_{t}$, is equal to the return of $\tilde{M}$, i.e. $f(\tilde{r}_{0},\dots,\tilde{r}_{T-1})$. To compute the immediate rewards $r_{t}$ of $M$ in a purely Markovian manner, we need access to information about the previous rewards of the trajectory. This can be achieved by extending the state space with $h_{t}$, which preserves all necessary information about the reward history. The function $u$ in Equation 4 updates this information at each time step. Finally, the Kronecker deltas in Equation 5 ‘pick’ all the possible $\tilde{r}_{t}$ that result in the same $h_{t+1}$ and $r_{t}$. For example, for $f(\tilde{r}_{0},\dots,\tilde{r}_{t})=\min(\tilde{r}_{0},\dots,\tilde{r}_{t})$, we can choose $r_{0}=h_{1}=\tilde{r}_{0}$ followed by $h_{t+1}=u(h_{t},\tilde{r}_{t})=\min(h_{t},\tilde{r}_{t})$, and $r_{t}=\rho(h_{t},\tilde{r}_{t})=\min(0,\tilde{r}_{t}-h_{t})$. More examples are shown in Table 1.
Due to Definition 1, there is a mapping between trajectories of NCMDPs and corresponding MDPs, which we will exploit to use the same policy for both decision processes:
Definition 2
A trajectory $\mathcal{\tilde{T}}=(\tilde{s}_{0},a_{0},\tilde{r}_{0},\dots)$ of an NCMDP $\tilde{M}$ is mapped onto a trajectory $\mathcal{T}=\mathrm{map}\left(\mathcal{\tilde{T}}\right)=(s_{0},a_{0},r_{0},\dots)$ of a corresponding MDP $M$ by calculating $r_{t}$ according to Equation 3, setting $s_{t}=(\tilde{s}_{t},h_{t})$ with $h_{t}$ calculated according to Equation 4, and keeping the actions $a_{t}$ the same.
Note that multiple $\mathcal{\tilde{T}}$ potentially map to the same $\mathcal{T}$ since $\rho$ and $u$, and, therefore, the map operation are not necessarily injective. We can now state the main result of this manuscript:
$\bm{f(\tilde{r}_{0}},\dots,\bm{\tilde{r}_{t})}$ Additional state information $\bm{h_{t}}$ Adapted reward $\bm{r_{t}}$
$\max(\tilde{r}_{0},\dots,\tilde{r}_{t})$ $h_{0}=-\infty$, $h_{t+1}=\max(h_{t},\tilde{r}_{t})$
$r_{0}=\tilde{r}_{0}$
$r_{t}=\max(0,\tilde{r}_{t}-h_{t}),t>0$
$\min(\tilde{r}_{0},\dots,\tilde{r}_{t})$ $h_{0}=\infty$, $h_{t+1}=\min(h_{t},\tilde{r}_{t})$
$r_{0}=\tilde{r}_{0}$
$r_{t}=\min(0,\tilde{r}_{t}-h_{t}),t>0$
Sharpe ratio $\frac{\mathrm{MEAN}(\tilde{r}_{0},\dots,\tilde{r}_{t})}{\mathrm{STD}(\tilde{r}_{0},\dots,\tilde{r}_{t})}$ $h_{0}=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}$, $h_{t+1}=\begin{bmatrix}\frac{h_{t}^{(2)}}{h_{t}^{(2)}+1}h_{t}^{(0)}+\frac{1}{h_{t}^{(2)}+1}\tilde{r}_{t}\\ \frac{h_{t}^{(2)}}{h_{t}^{(2)}+1}h_{t}^{(1)}+\frac{1}{h_{t}^{(2)}+1}\tilde{r}_{t}^{2}\\ \ h_{t}^{(2)}+1\end{bmatrix}$
$r_{t}=\frac{h_{t+1}^{(0)}}{\sqrt{h_{t+1}^{(1)}-h_{t+1}^{(0)^{2}}}}$
$-\frac{h_{t}^{(0)}}{\sqrt{h_{t}^{(1)}-h_{t}^{(0)^{2}}}}$
$\underset{k\in[-1,t]}{\max}\sum_{i=0}^{k}\tilde{r}_{i}$ $h_{0}=0$, $h_{t+1}=\max(0,h_{t}-\tilde{r}_{t})$ $r_{t}$ = $\max(0,\tilde{r}_{t}-h_{t})$
$\tilde{r}_{0}\tilde{r}_{1}\dots\tilde{r}_{t}$ $h_{0}=1$, $h_{t+1}=\tilde{r}_{t}h_{t}$
$r_{0}=\tilde{r}_{0},$
$r_{t}=h_{t+1}-h_{t},t>0$
Harmonic mean $\frac{1}{\frac{1}{\tilde{r}_{0}}+\dots+\frac{1}{\tilde{r}_{t}}}$ $h_{0}=0,h_{t+1}=h_{t}+\frac{1}{\tilde{r}_{t}}$
$r_{0}=\tilde{r}_{0},$
$r_{t}=\frac{1}{h_{t+1}}-\frac{1}{h_{t}},t>0$
$\delta^{t}\sum_{k=0}^{t}\tilde{r}_{k}$,
$\delta\in(0,1)$
$h_{0}=\begin{bmatrix}0\\ 0\end{bmatrix}$, $h_{t+1}=\begin{bmatrix}h_{t}^{(0)}+\tilde{r}_{t}\\ h_{t}^{(1)}+1\end{bmatrix}$
$r_{t}=\delta^{h_{t+1}^{(1)}}h_{t+1}^{(0)}$
$-\delta^{h_{t}^{(1)}}h_{t}^{(0)}$
$\frac{1}{t+1}\sum_{k=0}^{t}\tilde{r}_{k}$ $h_{0}=\begin{bmatrix}0\\ 0\end{bmatrix}$, $h_{t+1}=\begin{bmatrix}h_{t}^{(0)}+\tilde{r}_{t}\\ h_{t}^{(1)}+1\end{bmatrix}$
$r_{t}=\frac{1}{h_{t+1}^{(1)}}h_{t+1}^{(0)}$
$-\frac{1}{h_{t}^{(1)}}h_{t}^{(0)}$
@@
Theorem 1
Consider an NCMDP $\tilde{M}$ and a corresponding MDP $M$, both defined as in Definition 1. Then, given a policy $\pi(a_{t}|s_{t})$ of $M$, the expected return of $M$ is equal to the expected return of $\tilde{M}$, i.e.
$\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}r_{t}\right]=\mathbb{E}_{\pi}\left[f(\tilde{r}_{0},\dots,\tilde{r}_{T-1})\right],$ (6)
if $\pi(a_{t}|s_{t})$ is used to interact with $\tilde{M}$ as follows: At each time step $t$, map the current trajectory $\mathcal{\tilde{T}}$ of $\tilde{M}$ to a trajectory $\mathcal{T}=\mathrm{map}\left(\mathcal{\tilde{T}}\right)$ of $M$ as described in Definition 2, thereby finding $s_{t}$. Then choose an action according to $\pi(a_{t}|s_{t})$.
For a proof of this theorem see Appendix A. Since $M$ is a standard MDP, we can find its optimal policy using standard methods and are guaranteed that it will also be the ideal policy of $\tilde{M}$. More generally, also non-ideal policies will yield the same return for both decision processes. By mapping the NCMDP to an MDP, all guarantees available for MDP solvers, e.g. convergence proofs for Q-learning (Watkins and Dayan, 1992), or the policy gradient theorem for policy-based methods (Sutton et al., 1999), directly apply also to NCMDPs.
@@
We demonstrate the need for the additional state information $h_{t}$ in the two-step NCMDP depicted in Figure 2, which captures the main conceptual difficulties in solving NCMDPs: A non-cumulative objective in a stochastic environment. Here, the goal is to maximize the expected minimum of the rewards, i.e. $\min(\tilde{r}_{0},\tilde{r}_{1})$. If $\tilde{r}_{0}=1$ in the first step, the ideal policy is to choose $a_{1}$ in the second step, while if $\tilde{r}_{0}=-1$ the agent should choose $a_{0}$. Therefore, the choice of the ideal action depends on information contained in the past rewards, which is captured by $h_{1}=\tilde{r}_{0}$. We show the MDP constructed from the NCMDP as described above in the Supplement in Figure A6 and its state-action value function found by value iteration in Table A4. Our method finds the optimal policy, while previous methods for solving NCMDPs fail even in this simple example because they neglect the extra state information (see Table A5).
The question remains how to find the functions $u$ and $\rho$ given a new objective $f$. There is always at least one MDP corresponding to a given NCMDP since we can take $h_{t}=[\tilde{r}_{0},\dots,\tilde{r}_{t-1}]$ with $u(h_{t},\tilde{r}_{t})=[h_{t},\tilde{r}_{t}]$ and $\rho(h_{t},\tilde{r}_{t})=f([h_{t},\tilde{r}_{t}])-f(h_{t})$. However, this is not ideal because it leads to a state size that grows linearly with the trajectory length. In Section B.1, we provide a necessary and sufficient condition for objectives $f$ with constant size additional state information $h_{t}$. Moreover, we also provide a sufficient, but not necessary, condition that is easier to verify and covers a large class of functions, e.g. all functions in Table 1. For functions in this class, we also provide an explicit construction of $u$, $\rho$, and $h_{t}$ (see Section B.2). In practice, we have empirically observed that a simple analytical consideration leads to $h_{t}$ of a small and constant size for each of the objectives $f$ considered in this manuscript, which is desirable for efficient learning and integration with standard function estimators such as neural networks.
Limitation: Our method may result in a significantly increased state space, which might pose challenges for tabular reinforcement learning methods and dynamic programming techniques. However, deep reinforcement learning techniques are well equipped to address these challenges, as they can effectively handle exponentially large and continuous state spaces.
From an implementation perspective, our scheme of mapping NCMDPs to MDPs requires minimal effort, since we can treat both the NCMDP and the used MDP solver as black boxes by simply adding a layer between them, as shown in Figure 1. In addition, our treatment facilitates online learning and does not require any computationally expensive preprocessing of the NCMDP. This opens the door for researchers with specific domain knowledge, who are not necessarily experts in reinforcement learning, to use standard libraries such as stable-baselines3 (Raffin et al., 2021) to solve their non-cumulative problems. Conversely, it allows reinforcement learning experts to quickly tackle existing environments using our method.